OpenStudy (anonymous):

Evaluate the expression and write the result in the form a + bi. (9 + 3i)(3 − i) over/ 3 + i

OpenStudy (anonymous):

\[\frac{(9+3i)(3-i)}{3+i}\times\frac{3-i}{3-i}\]

OpenStudy (anonymous):

I got that far

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

\[\frac{(27-9i+9i-3i^2)*(3-i)}{9-i^2}\]

OpenStudy (anonymous):

\[\frac{(27-(3*-1)*(3-i)}{9-(-1)}\]

OpenStudy (anonymous):

\[\frac{(27+3)*(3-i)}{10}=\frac{90-30i}{10}\]

myininaya (myininaya):

oops i forgot to click post lol

myininaya (myininaya):

i had something different

OpenStudy (anonymous):

\[9-3i\]

myininaya (myininaya):

\[\frac{(9+3i)(3-i)}{3+i}=\frac{27-9i+9i-3i^2}{3+i}=\frac{27-3(-1)}{3+i}=\frac{30}{3+i}\] \[=\frac{30}{3+i}*\frac{3-i}{3-i}=\frac{90-30i}{9-i^2}=\frac{90-30i}{9-(-1)}=\frac{90-30i}{10}\] no you are right

myininaya (myininaya):

=9-3i

myininaya (myininaya):

i couldn't remember what i had before i forgot to click post lol

OpenStudy (anonymous):

ty